Sun Brite has a new pair of sunglasses it is evaluating. The company expects to sell 6,900 pairs of sunglasses at a price of $164 each and a variable cost of $116 each. The equipment necessary for the project will cost $360,000 and will be depreciated on a straight-line basis over the 9-year life of the project. Fixed costs are $300,000 per year and the tax rate is 35 percent. How sensitive is the operating cash flow to a $1 increase in variable costs per pairs of sunglasses?
Answer #1

Answer:

Annual depreciation = 360000 / 9 = $40,000

Annual operating cash flow = ((Sale price - Variable cost) * Units sold - Fixed cost) * (1 - Tax rate) + Depreciation tax shield

= ((164 - 116) * 6900 -300000) * (1 - 35%) + 40000 * 35%

= $34,280

Sensitivity of operating cash flow to a $1 increase in variable costs per pairs of sunglasses

Let us assume variable cost per unit increases by 1 to (116 +1=) $117

ΔOCF / ΔVC = (existing variable cost - new variable cost) *units sold * (1 - Tax rate)

= (116 - 117) * 6900 *(1 - 35%)

= - $4485

ΔOCF / ΔVC = - $4,485 or ($4485)

Sensitivity of operating cash flow to a $1 increase in variable costs per pairs of sunglasses = - $4,485 or ($4485)